8 + 4/x = 9 1/3 need help with some of the steps. Thank you for any help
Answer by stanbon(75887)

8 + 4/x = 9 1/3
----
8 + (4/x) = 28/3
--------
Multiply thru by 3x to get:
24x + 12 = 28x
4x = 12
x = 3
--------------
Cheers,
Stan H.
==============